OpenStudy (anonymous):

What are the foci of the hyperbola given by the equation 16x2 - 25y2 - 64x - 50y - 361 = 0?

OpenStudy (anonymous):

so isolate your x's and y's first so that you get.... \[(16x ^{2}-64x) - (25y ^{2}+50y) = 361\] then you do some factoring out so that you get.... \[(16(x ^{2}-4x)) - (25(y ^{2}+2y)) = 361\] now you do the squared rule thing to get... \[(16(x ^{2}-4x+4-4)) - (25(y ^{2}+2y+1-1)) = 361\] \[16(x-2)^{2}-64 - 25(y+1)^{2}+25 = 361\] then you just consolidate the x's and y's again.... \[16(x-2)^{2} - 25(y+1)^{2} = 400\] so now your center is (2, -1) now divide by 400 on each side to get.... \[\frac{ (x-2)^{2} }{ 25 } - \frac{ (y+1)^{2} }{ 16 } = 1\] from there you should be able to get your foci